A company produces​ 1,000 packages of cat food per month. The sales price is​ $4.00 per pack. Variable cost is​ $1.60 per​ unit, and fixed costs are​ $1,800 per month. Management is considering adding a vitamin supplement to improve the value of the product. The variable cost will increase from​ $1.60 to​ $1.80 per​ unit, and fixed costs will increase by​ 10%. The company will price the new product at​ $8 per pack. How will this affect operating​ income?
Answer #1
Current Proposed
Sales (1000*4)=4000 (1000*8)=8000
Variable costs (1000*1.6)=1600 (1000*1.8)=1800
Contribution margin 2400 6200
Fixed costs 1800 (1800*1.1)=1980
Operating income 600 4220

Hence increase in operating income=4220-600

=$3620